I would like to know how to solve this equation,(5/a-5)-1=(a/a-5).
Thanks to any help thats given.

Answer by lwsshak3(11628)

I would like to know how to solve this equation,(5/a-5)-1=(a/a-5).
Thanks to any help thats given.
**
(5/a-5)-1=(a/a-5)
LCD:(a-5)
5-a+5=a
2a=10
a=5
no solution because x≠5
if x=5,denominators=0, and the expression becomes undefined